\underline{Lectures on Motives}, (J. P. Murre), 9-16 May, 2005, Ankara.

Some selected References for Motives:

References

[1] A. Grothendieck, Standard conjectures on algebraic cycles, in Bombay Coll. Alg. Geom. 1968, Oxford Univ. Press, 1969.

[2] Y. Manin, Correspondences, motives and monoidal transformations, Math. USSR Sbornik 6, 1968.

[3] S. L. Kleiman, Motives in algebraic geometry, Conference Oslo, 1970.

[4] A. J. Scholl, Classical Motives, in Seattle Conf. on Motives 1991, AMS Proc. Sypos. P. Math. 55, 1994.

[5] J. P. Murre, Lectures on Motives, in Trancendental Aspects of Algebraic Cycles, Summer School Grenoble 2001, London Math. Soc. Lecture Notes, vol. 313, Cambridge Univ. Press, 2004.

[6] Y. André, Une introduction aux motifs, Panoramas et Synthèses, vol. 17, 2004, Soc. Math. de France.